OpenStudy (anonymous):

help!!!!!!!!!!!! step by step how do you do -4(2p+2)-44=2(3p+2)

OpenStudy (anonymous):

Take the LHS of the equation -4(2p+2)-44 = -4(2p) +(-4)2 - 44 = -8p-8-44=-8p-52

OpenStudy (anonymous):

Take now the RHS of the equation 2(3p+2)=2(3p)+2(2)=6p+4

OpenStudy (anonymous):

Now LHS=RHS -8p-52=6p+4 (add 8p on both sides of the equation) -8p+8p-52=6p+8p+4 -52=14p+4 (subtract 4 from both sides) -52-4=14p (divide by 14) -56/14=p p=4